# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### v2

@@ Abstract
Table-based reasoning has shown remarkable progress in combining deep models with discrete reasoning, which requires reasoning over both free-form natural language (NL) questions and semi-structured tabular data. However, previous table reasoning solutions only consider small-sized tables and exhibit limitations in handling larger tables. In addition, most existing methods struggle to reason over complex questions since they lack essential information or they are scattered in different places. To alleviate these challenges, we propose TAP4LLM as a versatile pre-processing toolbox to generate table prompts through (1) table sampling, (2) table augmentation, and (3) table packing while balancing the token allocation trade-off. In each module, we collect and design several common methods for usage in various scenarios (e.g., speed over accuracy). We also provide a comprehensive evaluation on performance of each components inside TAP4LLM and show that our method improves LLMs’ reasoning capabilities in various tabular tasks and enhances the interaction between LLMs and tabular data by employing effective pre-processing. The code of this paper will be released at https://anonymous.4open.science/r/TableProvider-4CC3.
@@
[…]
Yuan Sui1thanks: The contributions by Yuan Sui, Jiaru Zou, and Xinyi He have been conducted and completed during their internships at Microsoft Research Asia, Beijing, China., Jiaru Zou2***, Mengyu Zhou3thanks: Corresponding author., Xinyi He4***, Lun Du3, Shi Han3, Dongmei Zhang3 1 National University of Singapore 2 University of Illinois Urbana-Champaign 3 Microsoft 4 Xi’an Jiaotong University yuan.sui@u.nus.edu, jiaruz2@illinois.edu, hxyhxy@stu.xjtu.edu.cn {mezho, lun.du, shihan, dongmeiz}@microsoft.com
@@ 1 Introduction
The extensive and complex characteristics of data are commonly represented in the format of structured data. Table is one of those fundamental and widely used semi-structured data types in relational databases, spreadsheet applications, and programming languages that handle data for various domains, including financial analysis Zhang et al. (2020); Li et al. (2022), risk management Babaev et al. (2019), healthcare analytics Vamathevan et al. (2019), etc. Reasoning over tabular data is a fundamental aspect of natural language understanding (NLU) and information retrieval (IR), and has several downstream tasks, such as Table-based Question Answering (TQA) Chen et al. (2020b); Iyyer et al. (2017); Ye et al. (2023a); Cheng et al. (2023), Table-based Fact Verification (TFV) Chen et al. (2020a); Xie et al. (2022); Günther et al. (2021), Table-to-Text Wang et al. (2021b), Text-to-SQL Yu et al. (2018), Column Type & Relation Classification Iida et al. (2021); Deng et al. (2020),etc.
Figure 1: Demonstration of the Three TAP4LLM Modules.
@@
(1) Table sampling: sample most relevant content. (2) Table augmentation: retrieve and add extra information. (3) Table packing: serialize the sampled table and augmented information into a string while controlling the token allocation.
Meanwhile, large language models (LLMs) are advancing in their ability to tackle challenges associated with natural language using in-context learning Cheng et al. (2023); Ye et al. (2023a); Gemmell and Dalton (2023), but the degree to which they understand tables and how to leverage LLMs to work with table-based data remain an open question Chen (2022); Gong et al. (2020). Our research aims to explore this question and improve how LLMs use and work with table-based data.
@@
First, which part of a table should be kept in the prompt? The full content of a table could be very long and noisy to be included in the prompt. Most LLMs have a limited input context window size (e.g., 4k~16k tokens) in which an overlong table cannot fit. For long tables that satisfy the length constraint, it can still lead to unnecessary computations (of LLM on long prompt) and quality regressions (generation interfered by noisy input) when placing irrelevant table content (w.r.t. the task or query) in the prompt. To address the challenge, some sampling methods were proposed in ad-hoc ways. For example, truncating the input tables to contain only the first 20+ rows and 8 columns Chen (2022), or filtering relevant rows based on $n$-gram overlap between them and the utterance Yin et al. (2020). To answer the question of which part to keep, we need a more systematic study of different grounding and sampling algorithms.
@@
Second, what additional/external knowledge could help LLMs better understand a table? The raw content of a table may contain ambiguous information (e.g., abbreviations, domain-specific terms, column type,etc.) that requires further interpretation and clarification. As a result, direct reasoning on the raw tables may lead to misinterpretation and bias in the LLMs’ outputs. To address this, some augmentation techniques were proposed to incorporate structured knowledge Sui et al. (2023); Xie et al. (2022), common sense knowledge Bian et al. (2023); Ogundare and Araya (2023); Shen et al. (2023); Guo et al. (2023), and analytical knowledge Jena et al. (2022); He et al. (2023) into pre-training and inference processes. For example, Jena et al. (2022) transforms existing tabular data to create diverse NL inference instances for better zero-shot performance. He et al. (2023) infers implicit metadata behind raw table content through field distribution and knowledge graph information. However, the techniques were proposed independently and there lack a comprehensive study that compares them and tries to combine them to provide useful and diverse knowledge and thoughts for LLMs.
@@
Third, how do we encode the table into a prompt? While sampling and grounding compress the table content, augmentation expands the prompt by adding more information. With a given token budget, one needs to find the balance to allocate available tokens between table content and augmented knowledge. Furthermore, the serialization format of the table also plays a critical role. It not only influences how well an LLM understands the table input (as discussed in our previous study  Sui et al. (2023)), but also determines the string length of the serialized table and the augmented information. For example, as discussed in Sui et al. (2023), table formats such as HTML Aghajanyan et al. (2021) or XML are better understood by GPT models, but they also lead to increased token consumption. To pack a table into the prompt, these problems should be addressed with trade-offs.
In this paper, we propose TAP4LLM (table provider for large language models) as a versatile pre-processing toolbox to generate table prompts in LLM reasoning. TAP4LLM addresses the above challenges with three corresponding modules (i) Table Sampling: Selecting a sub-table $T^{\prime}$ from the raw table $T$ based on the rules or semantics of the query or utterance $Q$; (ii) Table Augmentation: Enhancing $T^{\prime}$ by integrating relevant external knowledge, metadata, and attributes based on the raw Table $T$; and (iii) Table Packing: Packing the sampled table $T^{\prime}$ with the augmented knowledge into a sequence with a specified serialization format for LLMs while balancing the token allocation trade-off. In each module, we collect and design several common methods for usage in various scenarios (e.g., speed over accuracy). Across six distinct datasets, our findings demonstrate that TAP4LLM significantly improves accuracy by achieving an average enhancement of 6.02% through the table sampling module, 3.29% through the table augmentation module, and 1.38% through the table packing module. Collectively, TAP4LLM elevates accuracy by an average of 7.93% when compared to the direct input of raw tables into LLMs. Our exploration has led us to conclude that TAP4LLM enhances the interaction between LLMs and tabular data by employing effective pre-processing.
@@
Figure 2: TAP4LLM Framework for Tabular Data. Noted that the “table sync” refers to the application (such as Excel Copilot) keeps its table data in sync with the table manager. The table manager acts as an intermediary, managing the data that is either stored locally in a cache or accessed through a database connection. This syncing process is crucial for “interactive table reasoning” and for maintaining data integrity. The implication of this syncing process is further discussed in §F.2.
@@
  • We proposed a unified pre-processor to improve the effectiveness of LLMs in tabular reasoning tasks and enhance the interaction between LLMs and tabular data.
  • We conducted a comprehensive evaluation of each components and showed that TAP4LLM achieves an average of improved performance by 7.93%.
  • We formulated a complete usage guideline for our framework TAP4LLM. For different real world scenarios, We identify corresponding optimal combination of methods within each module. we also provide a trade-off map between performance metrics and token allocation for reference.
@@ 2 TAP4LLM: Table Provider for LLMs
  • Table Augmentation: Explicitly incorporate relevant external knowledge, metadata, and attributes about the original table $T$.
  • Table Packing: Control the token allocation for table sampling and table augmentation; Convert the structured table into a sequence (table serialization).
@@ 2.1 Table Sampling
In table sampling, a subset of top-ranked rows and columns is selected to form the sub-table. Specifically, given an original table $T$ with a distinct set of rows $R_{T}$, columns $C_{T}$, and a query $q$, the goal of table sampling is to produce a sub-table $T^{\prime}=T_{r,c}$, where $r\in\mathcal{P}(R_{T})$, $c\in\mathcal{P}(C_{T})$. Here $\mathcal{P}(X)$ denotes the power set of $X$, representing all possible subsets of $X$. The process can be formulated as $T^{\prime}=T_{r,c}=\text{select}(T,\text{rank}(f(T,q)))$. The $f(T,q)$ function represents each sampling method. For example, the query-based sampling (discussed in details below) calculates the similarity score as $f$ between the query $q$ and each row / column from $T$. The $rank()$ function sorts the rows and columns of $T$ based on sampling methods $f$ and outputs a ranked list. The $select()$ function then chooses the top-k rows and top-l columns from the ranked list to form the sub-table $T_{r,c}$. Specifically, we classify multiple variants for table sampling as three following categories:
@@ 2.1.1 Rule-based Sampling
Rule-based sampling refers to table sampling based on predefined criteria or rules. These methods follow the established patterns or criteria for data selection. We consider three common rule-based sampling methods as the baselines: (1) Random Sampling, (2) Evenly Sampling and (3) Content Snapshot & Synthetically Sampling. The detailed description can be found in Appendix A.
@@ 2.1.2 Embedding-based Sampling
Instead of adhering to strict rules or criteria in rule-based sampling, embedding-based methods leverage the semantic and contextual representation of each row and column. Specifically, let $T$ be a table where $R_{T}$ is the set of rows and $C_{T}$ is the set of columns. Let $E:R_{T}\cup C_{T}\rightarrow\mathbb{R}^{d}$ be an embedding function that maps each row or column to a $d$-dimensional vector by capturing its semantic content. By mapping each row or column to vectors, this method harnesses the power of spatial relationships within the embedding space to guide sampling decisions. We propose two variant methods as follows:
(1) Semantic-based Sampling: Semantic-based Sampling is a tailored approach emphasizing the semantics relevance of row / columns to the utterance. The process is exactly illustrated in Eq. 2.1. Noted that the default query-based sampling is the row-based method. We also study the significance of column grounding as shown in Table 1.
(2) Centroid-based Sampling: Compared to the semantic-based sampling, the goal of centroid-based sampling is to ensure the preservation of data diversity. We use K-Means MacQueen et al. (1967) to partition the set of embeddings into $n$ clusters $C_{n}$. For each cluster $C_{i}$, we select the top-$K$ rows or columns based on the closeness to centroid. Since the construction of columns in a table inherently presupposes each column should be diversified, in that case, column-based sampling may not be well-suited for centroid-based methods.
(3) Hybrid-approach: The Hybrid approach marries the specificity of semantic-based sampling with the broad representations of centroid-based sampling. Compared to the centroid-based sampling, the top-$K$ rows or columns ($r$) are selected based on a combination metric $h(r,c,u)$ measuring the directional distance to cluster centroid ($c$) and the semantic similarity to the utterance ($u$). A straightforward approach to combine these two measures is: $h(r,c,u)=\alpha(\frac{1}{1+D(r,c)})+\beta S(r,u)$, where $D(r,c)$ measures the directional distance (e.g., Euclidean distance) between selected rows or columns and cluster centroid in embedding space, and $S(r,u)$ measures the semantic similarity between rows / columns and the utterance. The weights $\alpha$ and $\beta$ allow for flexible prioritization between these two aspects, tailoring the sampling process to emphasize either contextual relevance or diversity within the sampled data. In our experiment, we set $\alpha=0.3$ and $\beta=0.7$.
@@ 2.1.3 LLM-based Sampling
LLMs have been confirmed as effective decomposers for tabular reasoning Ye et al. (2023a). They employ a powerful LLM to facilitate sub-table extraction by predicting the indexes of rows and columns. However, reliance on LLMs for such pre-processing inevitably escalates computational costs and budgets. Moreover, using LLMs to predict the index still comes with challenges like token budget and noisy information and still requires table pre-processing. These issues inevitably transform the original table pre-processing task into a loop task. Despite this method not being ideally suited to our scope, we still consider it a strong baseline, albeit at the expense of time.
@@ 2.2 Table Augmentation
In table augmentation, we mainly consider the following three categories: (See the full knowledge aspects used in TAP4LLM from Table 6.)
@@ 2.2.1 Metadata-based Augmentation
Tabular data analysis have been evaluated to rely on accurate understanding of field semantics and further finding common patterns in daily analysis He et al. (2023). These kind of metadata in contrast to the raw tabular input which does not directly provide this information. Following AnaMeta He et al. (2023) using a range of knowledge-fusion models for metadata inference, we consider the following metadata-based augmentation types and leverage LLMs for zero-shot inference using metadata instruction as clues: Dimension / Measure, Semantic Field Type, Table Size, Statistics Feature, Header Hierarchy. The detailed description for each type can be found in Appendix C.

### v3

@@ Abstract
Table reasoning tasks have shown remarkable progress with the development of large language models (LLMs), which involve interpreting and drawing conclusions from tabular data based on natural language (NL) questions. Existing solutions mainly tested on smaller tables face scalability issues and struggle with complex queries due to incomplete or dispersed data across different table sections. To alleviate these challenges, we propose TAP4LLM as a versatile pre-processor suite for leveraging LLMs in table-based tasks effectively. It covers several distinct components: (1) table sampling to decompose large tables into manageable sub-tables based on query semantics, (2) table augmentation to enhance tables with additional knowledge from external sources or models, and (3) table packing & serialization to convert tables into various formats suitable for LLMs’ understanding. In each module, we design and compare several common methods under various usage scenarios, aiming to shed light on the best practices for leveraging LLMs for table-reasoning tasks. Our experiments show that our method improves LLMs’ reasoning capabilities in various tabular tasks and enhances the interaction between LLMs and tabular data by employing effective pre-processing.
@@ 1 Introduction
The extensive and complex characteristics of the data are commonly represented in the format of structured data. Table is one of those fundamental and widely used semi-structured data types in different areas, such as relational databases, spreadsheet applications, and programming languages that handle data for various domains, including financial analysis Zhang et al. (2020); Li et al. (2022), risk management Babaev et al. (2019), healthcare analytics Vamathevan et al. (2019), etc. Reasoning over tabular data has several important downstream tasks that are crucial to the field of natural language understanding (NLU) and information retrieval (IR), such as Table-based Question Answering (TQA) Chen et al. (2020b); Iyyer et al. (2017); Ye et al. (2023a); Cheng et al. (2023), Table-based Fact Verification (TFV) Chen et al. (2020a); Xie et al. (2022); Günther et al. (2021), Table-to-Text Wang et al. (2021b), Text-to-SQL Yu et al. (2018), Column Type & Relation Classification Iida et al. (2021); Deng et al. (2020), etc.
@@
Figure 1: Demo of TAP4LLM Modules. (1) Table sampling: sample most relevant content. (2) Table augmentation: retrieve and add extra / meta information. (3) Table packing: serialize the sampled table and augment information into a string while controlling the number of tokens.
Although there is remarkable progress on table reasoning tasks with the development of large language models (LLMs) Cheng et al. (2023); Ye et al. (2023a); Gemmell and Dalton (2023), the existing solutions are mainly tested on small tables and cannot reflect the real challenges of table reasoning. For example, there are usually issues on scaling to large tables or handling complex queries that require gathering data from various parts of a table. How well do LLMs understand tables and how to leverage LLMs to work with table data remains an open question Chen (2022); Gong et al. (2020). Our research aims to explore this question and shed light on the best practices to leverage LLMs for table-reasoning tasks. Several specific practical issues are faced when leveraging LLMs for table reasoning tasks as follows.
@@
First, which part of a table should be kept in the prompt? The full content of a table could be too lengthy and noisy to be included in the prompt. In addition, most LLMs have a limited input context window size in which an overlong table cannot fit. For large tables that satisfy the length constraint, it can still lead to unnecessary computations (of LLM on long prompt) and quality regressions (generation interfered from noisy input) when placing irrelevant table content (w.r.t. the task or query) in the prompt. To address the challenge, some sampling methods were proposed in ad-hoc ways. For example, truncating the input tables to contain only the first 20+ rows and 8 columns Chen (2022), or filtering relevant rows based on $n$-gram overlap between them and the utterance Yin et al. (2020). To answer the question of which part to keep, we conduct a systematic study of different grounding and sampling algorithms in Section 2.1, and the experiments and findings can be found in Section 3.2.
@@
Second, what additional/external knowledge could help LLMs better understand a table? The raw content of a table may contain ambiguous information (e.g., abbreviations, domain-specific terms, column type, etc.) that requires further interpretation and clarification. As a result, direct reasoning with the raw table may lead to misinterpretation and hallucination by LLMs. To address this, some augmentation techniques were proposed to incorporate structured knowledge Sui et al. (2023); Xie et al. (2022), common sense knowledge Bian et al. (2023); Ogundare and Araya (2023); Shen et al. (2023); Guo et al. (2023), and analytical knowledge Jena et al. (2022); He et al. (2023) into training and inference processes. For example, Jena et al. (2022) transforms existing tabular data to create diverse NL inference instances for better zero-shot performance. AnaMeta He et al. (2023) infers implicit metadata behind raw table contents through field distribution and knowledge graph information. However, the techniques were proposed independently and there lack a comprehensive study that compares them and attempts to combine them to provide useful and diverse knowledge and thoughts for LLMs. We will discuss several augmentation methods in Section 2.2 and their corresponding experiments and findings can be found in Section 3.3.
Moreover, table augmentation plays a crucial role in avoiding LLMs to partitionally comprehend the semantics and distribution of the whole table solely based on table sampling (which may remove some essential rows/columns due to the limitations of the methods). It leverages the summarization, statistics, and metadata information derived from the whole table to represent high-level information, to compromise the trade-off with table sampling, which will intuitively decrease information entropy. We will discuss this essential trade-off in Section 3.5 and the experiments and findings can be found in Figure 3.
@@
Third, how do we encode the table into a prompt? While sampling and grounding compress the table content, augmentation expands the prompt by adding more information. With a given token budget, one needs to find the balance to allocate available tokens between table content and augmented knowledge. Furthermore, the serialization format of the table also plays a critical role. It not only influences how well an LLM understands the table input Sui et al. (2023), but also determines the string length of the serialized table and the augmented information. For example, as discussed in Sui et al. (2023), table formats such as HTML Aghajanyan et al. (2021) or XML are better understood by GPT models, but they also lead to increased token consumption. To pack a table into the prompt, these problems should be addressed with trade-offs (see Section 3.5).
@@
Figure 2: TAP4LLM Framework for Tabular Data. Note that “table sync” refers to the application (such as Excel Copilot) keeping its table data in sync with the table manager. The table manager acts as an intermediary, managing the data that is either stored locally in a cache or accessed through a database connection. This sync process is crucial for “interactive table reasoning” and for maintaining data integrity. The implications of this syncing process are further discussed in §F.2.
To mitigate the aforementioned challenges, we propose TAP4LLM (TAble Provider for Large Language Models) as a versatile pre-processor suite for LLMs in table-based tasks. Specifically, there are three essential modules: (i) Table Sampling: decomposing large tables into manageable sub-tables based on query semantics; (ii) Table Augmentation: enhancing tables with additional knowledge from external sources or symbolic models; and (iii) Table Packing: convert tables into various formats (e.g., HTML, XML, Markdown, etc) suitable for LLMs’ understanding while balancing the token allocation trade-off as well.
In each module, we design and compare new and existing methods for various scenarios across six distinct datasets. Through experiments in Section 3, we find that: (1) When using LLMs to process tables, it is more effective to concentrate on key rows and columns rather than overloading with extraneous data. For tasks that require high accuracy, semantic-based sampling typically enhances performance, whereas for tasks prioritizing low latency and minimal computational resources, rule-based sampling may be more suitable. (2) Integrating external knowledge of the tables can consistently improve the performance of table reasoning tasks by reducing hallucinations and factual inaccuracies in LLMs and improving general comprehension and analysis of tabular data. (3) A balanced distribution of tokens between table content and augmented information can help improve overall performance; We find a performance-optimized ratio (close to 5:5 or 4:6 between table content and augmentation tokens) often achieves the best performance across different settings.
@@
  • We propose TAP4LLM framework to improve the effectiveness of LLMs on tabular reasoning tasks by better sampling, augmentation and packing tables into input prompts.
  • A comprehensive evaluation of each component is conducted. On average TAP4LLM achieves a 7.93% performance improvement.
  • We summarize a table prompting guideline. For different real-world scenarios, we identify the corresponding optimal combination / setting of methods within each module.
@@ 2 TAP4LLM: Table Provider for LLMs
  • Table Augmentation: Incorporate relevant external knowledge, metadata, and attributes about the original table $T$ explicitly.
  • Table Packing: Convert tables into various formats suitable for LLMs’ understanding while control the token allocation for table sampling and augmentation.
@@ 2.1 Table Sampling
In table sampling, a subset of top-ranked rows and columns is selected to form a sub-table. Specifically, given an original table $T$ with a distinct set of rows $R_{T}$, columns $C_{T}$, and a query $q$, the goal of table sampling is to produce a sub-table $T^{\prime}=T_{r,c}$, where $r\in\mathcal{P}(R_{T})$, $c\in\mathcal{P}(C_{T})$. Here $\mathcal{P}(X)$ denotes the power set of $X$, representing all possible subsets of $X$. The process can be formulated as
$T^{\prime}=T_{r,c}=select(T,rank(f(T,q)))$ (1)
The $f(T,q)$ function represents each sampling method. For example, the query-based sampling (discussed in detail below) calculates the similarity score as $f$ between the query $q$ and each row/column from $T$. The $rank()$ function sorts the rows and columns of $T$ based on sampling methods $f$ and outputs a ranked list. The $select()$ function then chooses the top-k rows and top-l columns from the ranked list to form the sub-table $T_{r,c}$. Specifically, we classify multiple variants for table sampling as three following categories:
@@ 2.1.1 Rule-based Sampling
Rule-based sampling refers to table sampling based on predefined criteria or rules. These methods follow the established patterns or criteria for data selection. Specifically, we consider three common rule-based sampling methods incoporated into our table sampling module: (1) Random Sampling, (2) Evenly Sampling, and (3) Content Snapshot & Synthetically Sampling. The detailed description can be found in Appendix A.
@@ 2.1.2 Embedding-based Sampling
Instead of adhering to strict rules or criteria in rule-based sampling, embedding-based methods leverage the semantic and contextual representation of each row and column. Specifically, let $T$ be a table where $R_{T}$ is the set of rows and $C_{T}$ is the set of columns. Let $E:R_{T}\cup C_{T}\rightarrow\mathbb{R}^{d}$ be an embedding function that maps each row or column to a $d$-dimensional vector by capturing its semantic content. By mapping each row or column to vectors, this method harnesses the power of spatial relationships within the embedding space to guide sampling decisions.
Here, we propose three variant methods as shown below: (1) Semantic-based Sampling: Semantic-based Sampling is a tailored approach emphasizing the semantics relevance of row/columns to the utterance. The process is exactly illustrated in Eq. 1. Note that the default query-based sampling is the row-based method. In our experiments, we also study the column grounding shown in Table 1. (2) Centroid-based Sampling: The goal of centroid-based sampling is to ensure the preservation of data diversity. We use K-Means MacQueen et al. (1967) to partition the set of embeddings into $n$ clusters $C_{n}$. For each cluster $C_{i}$, we select the top-$K$ rows or columns based on the closeness to the centroid. (3) Hybrid-approach: The Hybrid approach marries the specificity of semantic-based sampling with the broad representations of centroid-based sampling. Specifically, the top-$K$ rows or columns are selected based on a combination metric $h(r,c,u)$ measuring the directional distance to cluster centroid $c$ and the semantic similarity to the utterance $u$. We formulate the measuring metric as:
$h(r,c,u)=\alpha(\frac{1}{1+D(r,c)})+\beta S(r,u)$ (2)
where $D(r,c)$ measures the directional distance (e.g., Euclidean distance) between selected rows or columns and cluster centroid in embedding space, and $S(r,u)$ measures the semantic similarity between rows/columns and the utterance. The weights $\alpha$ and $\beta$ provide flexibility in prioritizing between contextual relevance and diversity. Without further specification, we set $\alpha=0.3$ and $\beta=0.7$ in our experiments.
@@ 2.1.3 LLM-based Sampling
LLMs have proven effective in tabular reasoning Ye et al. (2023a), utilizing their capabilities to predict row and column indices for efficient sub-table extraction. However, relying on LLMs for pre-processing significantly increases computational costs. Additionally, using LLMs to predict indices introduces challenges such as token limitations, noisy information, and the need for further table pre-processing, turning the task into a recursive loop. Despite this method not being ideally suited to our scope, we still consider it a strong baseline, albeit at the expense of time.
@@ 2.2 Table Augmentation
Table augmentation enhances LLM reasoning by adding extra knowledge to the input table and query. In the table augmentation module, we categorize different knowledge aspects into three main categories: Metadata-based, Retrieval-based, and Self-consistency-based (See Table 7 in Appendix). We will describe each category in detail below.
@@ 2.2.1 Metadata-based Augmentation
Tabular data analysis relies on accurately understanding field semantics and identifying common patterns in everyday analysis He et al. (2023). Following AnaMeta He et al. (2023) using a range of knowledge-fusion language models for metadata inference, we consider five main metadata-based augmentation types and leverage LLMs for zero-shot inference using metadata instruction as clues: Dimension / Measure, Semantic Field Type, Table Size, Statistics Feature, Header Hierarchy. The detailed description for each meta-data augmentation type can be found in Appendix C.